OpenStudy (anonymous):

|[x+2]| when x=-2,1 [|x+2|] when x=-2,1 plz help :(

OpenStudy (anonymous):

x=-2.1 is it true?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

x=2 and x=1 ya?

OpenStudy (anonymous):

i mean one time x=2 and the other time is x=1

OpenStudy (anonymous):

x=-2

OpenStudy (anonymous):

x=-2

OpenStudy (anonymous):

got it:)

OpenStudy (anonymous):

x=-2 so x+2=0 and [x+2] = |[x+2]| = 0

OpenStudy (anonymous):

okey

OpenStudy (anonymous):

x=1; x+2=3 and [x+2]=|[x+2]|=3 is it ok?

OpenStudy (anonymous):

yeah i did the same

OpenStudy (anonymous):

is there any problem else?

OpenStudy (anonymous):

how about the second ?

OpenStudy (anonymous):

[|x+2|]

OpenStudy (anonymous):

oh. x=-2 so |x+2|=0 and [|x+2|] =0

OpenStudy (anonymous):

x=1; |x+2| =3 and [|x+2|]=3 ok?

OpenStudy (anonymous):

thank`s alot hamed

OpenStudy (anonymous):

you are welcome:)

OpenStudy (anonymous):

after long time when i want to solved math , it `s come funny to me

OpenStudy (anonymous):

no. u only should more practice:)

OpenStudy (anonymous):

i solved it

OpenStudy (anonymous):

just in doubt

OpenStudy (anonymous):

it's natural for u at first. don't worry:)

OpenStudy (anonymous):

it seams you are iranian , aren`t u?

OpenStudy (anonymous):

yes:)

OpenStudy (anonymous):

good

OpenStudy (anonymous):

glad to meet u dear Pasha:)

OpenStudy (anonymous):

cya later

OpenStudy (anonymous):

same here,hamed

OpenStudy (anonymous):

hope 2

OpenStudy (anonymous):

khodahafez

OpenStudy (anonymous):

khoda negahdar:)